# Thermal Time and Irreversibility from Non-Commuting Observables in Accelerated Quantum Systems

### 09 Apr 2026 v1

@@ 6 Discussion
We have identified a minimal mechanism through which temporal ordering becomes operationally meaningful in relativistic quantum systems. The key ingredients are the KMS structure of the underlying state and the non-commutativity of the observables through which the detector couples to the field.
At the dynamical level, the ordering of interactions is encoded in the reduced detector state. Distinct interaction protocols lead to different final states whenever non-commutativity produces a nonvanishing contribution $\Delta\rho_{D}$. In the presence of a KMS state, this asymmetry acquires a universal thermal structure controlled by the local temperature. This establishes that temporal ordering can be reflected directly in the state of a localized quantum system.
The quantitative content of this asymmetry is captured by relative entropy. In the minimal two-level model, the relevant states form a family of Gibbs states with identical spectra and different generators. The relative entropy between them depends only on the dimensionless parameter $s=\beta\Delta$, linking the strength of the asymmetry to the local temperature scale.
@@
The comparison between the Bogoliubov–Kubo–Mori and Bures metrics clarifies the geometric structure underlying this result. In commuting directions the two metrics coincide, while in non-commuting directions they differ. This difference quantifies the separation between irreversible entropic cost and operational distinguishability, and is governed by the same thermal parameter.
Thermal time is realized in this framework through the modular generator $K=-\log\rho$ associated with a Gibbs state. In the accelerated setting, this generator is normalized by the local inverse temperature $\beta(\xi)$, determined by the Tolman relation and the Unruh effect. The dependence of the relative entropy and of the information-geometric quantities on $s=\beta\Delta$ makes explicit how irreversibility is controlled by this local thermal scale.
The detector model provides a concrete probe of these structures. The KMS condition fixes the analytic properties of the correlation functions, while non-commuting observables probe this structure through sequential interactions. The resulting effective Gibbs states encode the ordering dependence, and their relative entropy quantifies its irreversible content.
@@
The framework suggests several directions for further investigation. One extension is to consider detector systems with higher-dimensional internal Hilbert spaces, where the structure of non-commuting generators is richer. Another is to analyze other spacetime settings in which KMS states arise, such as stationary spacetimes with horizons, where nontrivial temperature profiles may lead to additional effects.
The results show that temporal ordering acquires physical meaning when it is encoded in the state of a system through the interplay of thermality and non-commutativity. In this setting, irreversibility is quantified by relative entropy, and its magnitude is set by the same thermal scale that governs the underlying KMS structure.
Appendix A Relative entropy for non-commuting Gibbs states
We derive the closed-form expression for the relative entropy between two Gibbs states generated by non-commuting operators. This computation provides the entropic measure used in the main text to quantify the mismatch between the corresponding non-commuting Gibbs structures.
@@ Appendix A Relative entropy for non-commuting Gibbs states
Using the identity
@@
The logarithm of the density matrix follows directly from its Gibbs form,
@@
The quantum relative entropy is defined as
@@
Substituting the expressions above, one finds
@@
The expectation values of Pauli operators in the state $\rho_{y}$ are
$\mathrm{Tr}(\rho_{y}\sigma_{y})=-\tanh s,\qquad\mathrm{Tr}(\rho_{y}\sigma_{x})=0.$
It follows that
@@
More generally, if one interpolates between the two generators by
$H_{\theta}=\Delta(\cos\theta\,\sigma_{x}+\sin\theta\,\sigma_{y}),$
the second-order expansion of relative entropy with respect to $\theta$ yields the Bogoliubov–Kubo–Mori metric used in Section 5.
The expression for $D(\rho_{y}\|\rho_{x})$ depends only on the dimensionless parameter $s=\beta\Delta$ and therefore encodes the dependence of the irreversible informational cost on the local temperature scale. In the context of the main text, this quantity provides the entropic measure of the mismatch between Gibbs states generated by non-commuting observables.
@@ Appendix B Perturbative derivation of the ordering asymmetry
For stationary trajectories, one may introduce the Fourier transform
@@
This shows explicitly how the KMS condition induces a thermal weighting of the symmetrized correlator that governs the ordering asymmetry.

### 27 Apr 2026 v2

@@ 6 Discussion
We have identified a minimal mechanism through which temporal ordering becomes operationally meaningful in relativistic quantum systems. The key ingredients are the KMS structure of the underlying state and the non-commutativity of the observables through which the system probes the field.
The mechanism can be summarized as follows. The KMS condition imposes a specific analytic structure on correlation functions, which translates into a frequency-dependent weighting of forward and backward processes. When the detector couples through non-commuting observables, this structure is reflected in the reduced detector state as a difference between distinct interaction protocols. The resulting asymmetry appears at the level of the density matrix and does not require coarse-graining or macroscopic limits.
In the minimal two-level model, the detector states associated with different internal observables are described by effective Gibbs distributions with identical spectra and different generators. The relative entropy between these states provides a natural measure of their difference. Its explicit form shows that the same parameter controlling the Unruh response also governs the strength of the ordering asymmetry.
@@
The comparison between the Bogoliubov–Kubo–Mori and Bures metrics clarifies the distinction between entropic and operational aspects of this effect. In commuting directions, these metrics coincide, reflecting the classical character of thermal fluctuations. In non-commuting directions, they differ, and this difference is controlled by the same parameter that determines the thermal response. The resulting structure links distinguishability, irreversibility, and local thermal scale in a unified way.
The analysis presented here is based on perturbation theory and localized detector models. It therefore inherits the assumptions of that framework, including the use of switching functions and the restriction to leading nontrivial order in the coupling. Extending the results beyond this regime would require a non-perturbative treatment of detector dynamics or an explicit construction within algebraic quantum field theory.
@@
The framework suggests several directions for further investigation. One natural extension is to consider more general detector systems, including higher-dimensional internal Hilbert spaces or continuous-variable models, where the structure of non-commuting observables is richer. Another direction is to study different spacetime settings in which KMS states arise, such as stationary spacetimes with horizons, where similar mechanisms may be present.
The relation between modular structure and information geometry may also be explored further. In particular, the role of relative entropy as a measure of modular mismatch suggests a possible link with other contexts in which modular Hamiltonians play a central role, such as quantum field theory in curved spacetime and quantum information theory.
The results show that KMS thermality, modular structure, and non-commutativity jointly determine when temporal ordering becomes observable and when it acquires a preferred direction. In this sense, the arrow of thermal time emerges as the direction along which relative entropy with respect to the KMS state increases.
A natural extension of this work is to analyze whether similar ordering asymmetries arise in other settings where KMS states are present, such as stationary spacetimes with horizons. It would also be of interest to explore non-perturbative formulations of the effect and its relation to modular Hamiltonians in algebraic quantum field theory.
Appendix A Derivation of relative entropy for non-commuting Gibbs states
In this appendix we derive the closed-form expression for the relative entropy between the two effective Gibbs states introduced in Section 5.
@@ Appendix A Derivation of relative entropy for non-commuting Gibbs states
Using the identity for Pauli matrices
@@
The logarithm of the density matrix is
@@
The relative entropy is defined as
@@
Substituting the expressions above gives
@@
Using
$\mathrm{Tr}(\rho_{y}\sigma_{y})=-\tanh s,\qquad\mathrm{Tr}(\rho_{y}\sigma_{x})=0,$
we obtain
@@ Appendix B Perturbative derivation of the ordering asymmetry
For stationary trajectories, one may introduce the Fourier representation
@@
This shows explicitly that the ordering asymmetry is governed by the symmetrized correlation function of the field, with a thermal weighting fixed by the KMS condition.